find the slope-interception equation of the line passing through the point (2, 1) with a slope of m = 3
would appreciate your help

Answer by ikleyn(52904)

An equation of a straight line in a coordinate plane which has the slope "m" and passes through the given point  P = (a,b)  is 


y - b = m*(x-a).               (1)          ( * * * MEMORIZE it ! * * * )


Now substitute the given data  m = 3,  a= 2,  b= 1  into equation  (1),  and you will get

y-1 = 3*(x-2).                 (2)


or, equivalently, 


y = 3x - 6 + 1,   which is   y = 3x - 5.


The last equation is in the  "slope intercept form", and it is your answer.

---------------
See the lesson
    - Equation for a straight line having a given slope and passing through a given point
in this site.